OpenStudy (anonymous):

C= 5/9(F-32) SOLVE FOR F

ganeshie8 (ganeshie8):

to solve the equation for F you need to isolate F

ganeshie8 (ganeshie8):

since the expression on right, is being divided by 9, perform the opposite operation - start with multiplying both sides of equal sign with 9

ganeshie8 (ganeshie8):

=> \(C = \frac{5}{9}(F-32) \) \(9*C = 9*\frac{5}{9}(F-32) \) \(9C = 5(F-32) \)

ganeshie8 (ganeshie8):

again repeat the same thing perform the opposite operation - since the right side is multiplied by 5, divide both sides with 5

ganeshie8 (ganeshie8):

\(\frac{9C}{5} = \frac{5}{5}(F-32)\) \(\frac{9C}{5} = F-32\)

ganeshie8 (ganeshie8):

Finally add 32 both sides to isolate the F completely !

OpenStudy (anonymous):

C=5/9(F-32) 9C=5(F-32) 9C/5=F-32 9C+32/5=F

ganeshie8 (ganeshie8):

\(\frac{9C}{5}+32 = F\) or \(F = \frac{9C}{5}+32\) you can leave the expression either way. both are same !!

OpenStudy (anonymous):

thank you!!!!!!!

ganeshie8 (ganeshie8):

yw :))